OpenStudy (anonymous):

you have these bnumbers : 7,7,7,7,1 (you cannot use 7 more then 4 times, and 1 more then 1 time, ) you have to get use, by using ummation, extraction, multiplication or division. :)

hero (hero):

What's the question?

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

whops, I forgot, your goal is to get 100

jhonyy9 (jhonyy9):

7*7=49+1=50*(7+7)/7=100

OpenStudy (anonymous):

You used 7 five times, but the limit is 4...

jhonyy9 (jhonyy9):

7 squared wann to be two 7 or just one ?

OpenStudy (anonymous):

7^7 = 2 7...